OpenStudy (anonymous):

How do you implicitly differentiate: x^3y^3-y=x

OpenStudy (anonymous):

take y to the other side then apply product rule in the left hand side

OpenStudy (anonymous):

gimme a sec i am working on it

OpenStudy (amistre64):

you aint gots to move anything, just differentiate

OpenStudy (anonymous):

haha amistre is correct

OpenStudy (amistre64):

:) it happens

OpenStudy (amistre64):

x^3y^3-y=x x'^3y^3+x^3y^'3-y'=x'

OpenStudy (amistre64):

eventually you factor out a y' and move it all to the other side tho

OpenStudy (anonymous):

d/dx(x^3y^3)-d/dx(y)=d/dx(x)

OpenStudy (anonymous):

For d/dx(x^3y^3) you use the Chain Rule?

OpenStudy (amistre64):

product rule, and some say chain rule for the y(x) part

OpenStudy (anonymous):

What are the steps please?

OpenStudy (anonymous):

\[x^3y^3-y=x\] is like \[x^3f(x)^3-f(x)=x\]

OpenStudy (anonymous):

for first term you use the product rule and the chain rule \[3x^2f^3(x)+x^33f^2(x)f'(x)\] or \[3x^2y^3+3x^3y^2y'\]

OpenStudy (anonymous):

so start with \[3x^2y^3+3x^3y^2y-y'=1\] then solve for \[y'\]

OpenStudy (anonymous):

rather \[3x^2y^3+3x^3y^2y'-y'=1\]

OpenStudy (anonymous):

So the answer is: \[y' = (1-3x^2y^3)/(3x^2-1)\]

OpenStudy (amistre64):

forgot a y^2 in the denom

OpenStudy (amistre64):

but other than that, yes

OpenStudy (anonymous):

So it's \[y′=(1−3x^2y^3)/(3x^2−y^2)\]

OpenStudy (amistre64):

\[y′=\frac{1−3x^2y^3}{3x^2y^2−1}\]

OpenStudy (anonymous):

I'm trying to figure out where I forgot the y^2..

OpenStudy (amistre64):

\[3x^2y^3+3x^3y^2y'-y'=1\] \[3x^3y^2y'-y'=1-3x^2y^3\] \[y'(3x^3y^2-1) =1-3x^2y^3\]

OpenStudy (anonymous):

Yes, the correct one is: y' = ( 1- 3x^2 y^3 ) / ( 3x^3 y^2 -1)